### v1

# VoroUDF: Meshing Unsigned Distance Fields with Voronoi Optimization

Abstract.
We present VoroUDF, an algorithm for reconstructing high-quality triangle meshes from Unsigned Distance Fields (UDFs). Our algorithm supports non-manifold geometry, sharp features, and open boundaries, without relying on error-prone inside/outside estimation, restrictive look-up tables nor topologically noisy optimization. Our Voronoi-based formulation combines a $L_{1}$ tangent minimization with feature-aware repulsion to robustly recover complex surface topology. It achieves significantly improved topological consistency and geometric fidelity compared to existing methods, while producing lightweight meshes suitable for downstream real-time and interactive applications.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
Figure 1. We propose an algorithm for reconstructing surfaces with non-manifold structures (highlighted in red), sharp features, and open boundaries from unsigned distance fields with gradient information.
1. Introduction
Figure 2. Unlike marching-cube-based methods (Guillard et al., 2022; Zhou et al., 2022; Ren et al., 2023; Stella et al., 2024) that rely on pseudo-sign estimation, double-covering-based methods (Hou et al., 2023; Chen et al., 2025) that produce double-layered meshes, and dual-contouring-based methods (Zhang et al., 2023) that introduce erroneous connectivity, our method is the only approach capable of reconstructing non-manifold structures (highlighted in red).
Unsigned Distance Fields (UDFs) have emerged as a powerful representation for shapes of arbitrary topology containing sharp and even non-manifold features. Their ability to represent thin, geometrically challenging self-intersecting structures has made them popular in artistic 3D modeling applications (see Figure 1), while their flexibility has increased their adoption in learning-based shape representations and reconstruction pipelines for real-world geometry. Reconstructing high-quality surface meshes with non-manifold structures, open boundaries, and sharp features from UDFs is critical for downstream applications such as texture mapping, physical simulation, animation, and geometric processing.
However, converting a UDF into an explicit surface mesh remains challenging. Without a consistent sign to differentiate inside from outside, classic grid-based iso-surfacing techniques (Lorensen and Cline, 1987; Ju et al., 2002) cannot be directly applied. Methods that combine sign estimation with Marching Cubes lookup tables (Guillard et al., 2022; Zhou et al., 2022; Ren et al., 2023; Stella et al., 2024) are highly unstable around thin regions and fail to reconstruct non-manifold structures (see Figure 1, second and third from left). Dual grid contouring approaches with purpose-made quadratic optimizations (Zhang et al., 2023) produce topological noise and undesirable non-manifold edges missing from the input (Figure 1, second from right). When non-manifold features are intentionally present in the input (e.g., Figure 1, left), algorithms based on contracting a surface mesh of a volumetric offset (Hou et al., 2023; Chen et al., 2025) cannot recover them by design (Figure 1, third from right).
In this paper, we introduce a Voronoi-based optimization framework for extracting isosurfaces with non-manifold structures, open boundaries, and sharp features from UDFs. In contrast to recent grid-based methods that rely on fixed spatial partitions, our Voronoi-based decomposition adapts to the distribution of discrete seeds in space, enabling a more flexible and geometry-aware partitioning. We iteratively solve a local $L_{1}$ tangent energy minimization and a global seed distribution problem to optimize seed positions such that they automatically align with complex surface features. Connectivity is directly derived from the Voronoi adjacency, resulting in cleaner and more topologically consistent meshes than those produced by grid-based methods (Figure 1, right).
Through extensive qualitative and quantitative evaluations on synthetic non-manifold shapes and real-world datasets (Koch et al., 2019; Zhu et al., 2020), we demonstrate that our method outperforms prior work in reconstructing non-manifold structures, open boundaries, and sharp features. Our algorithm will be most impactful in artistic and industrial design contexts, as is it able to recover the geometrically challenging, topologically degenerate shapes common in these. Critically, it does so without resorting to an excessively dense discretization, producing lightweight meshes that are directly portable to realtime and interactive applications.
2. Related Work
Figure 3. Illustration of why marching-cubes-based methods (top row), double-covering-based methods (middle row), and dual-contouring-based methods (bottom row) fail to preserve the non-manifoldness of the iso-surface.
UDFs can represent surfaces of arbitrary topology, making them particularly valuable in domains such as artistic modeling, manufacturing, and industrial design, where accurate recovery of complex surface structures is critical. Owing to this flexibility, UDFs have become a powerful tool across a wide range of scientific fields—from medical imaging (Jin et al., 2024; Sørensen et al., 2022) to geology (Zhang et al., 2026) and additive manufacturing (Brunton and Rmaileh, 2021). In computer graphics, their expressive capacity has been extensively utilized in geometric deep learning and generative 3D modeling (Xu et al., 2025; Chibane et al., 2020; Liu et al., 2023; Long et al., 2023; Zhao et al., 2021).
Despite their expressiveness, UDFs remain poorly suited for many downstream applications due to its implicit nature. Tasks such as finite element simulation, texture mapping, real-time rendering, and animation often require converting UDFs into explicit surface representations like polygonal meshes. This conversion is nontrivial: UDFs lack orientation and sign information, making direct mesh extraction challenging. As a result, robust UDF-to-mesh reconstruction remains a central problem in geometry processing. Several recent strategies have been explored to reconstruct a surface mesh from a UDF, each with important limitations.
Marching-cubes-based methods
One family of methods adapts grid-based isosurface extraction (e.g., Marching Cubes variants) to UDFs by introducing a pseudo-sign per grid cell (Guillard et al., 2022; Zhou et al., 2022; Ren et al., 2023). These approaches attempt to infer on which side of the surface each voxel corner lies (since all UDF values are positive) by analyzing the UDF’s gradients or other heuristics. Essentially, they locally treat the UDF as a signed field and then apply the Marching Cubes lookup to extract a mesh. This works only under a key assumption: each grid cell that contains the surface must lie entirely on one side of the surface’s medial axis. If a cell spans across the medial axis, the pseudo-sign becomes ill-defined and unstable (see Fig. 3 top row). The learning-based method NSDUDF (Stella et al., 2024) predicts pseudo-sign information from local views, however, it struggles to generalize and may fail to produce accurate results on previously unseen shapes or topologies, particularly in regions with non-manifold geometry.
Double-covering-based methods
Another approach is to temporarily convert the open or non-manifold surface into a closed manifold by using a double covering technique (Hou et al., 2023; Chen et al., 2025). This approach extracts a thin offset surface around the target shape, which forms a closed, orientable manifold, and then optimizes this double-layered surface onto the direction of isosurfaces. This yields a valid mesh even when the input surface is open. If the input surface is truly non-manifold, however, the method cannot produce a single unified mesh, as it instead retains two overlapping surface layers at the non-manifold regions (see Fig. 3 middle row).
Dual-contouring-based methods
To avoid relying on a global pseudo-sign, researchers have also proposed dual-contouring-based methods for UDF meshing. The learning-based method NDC (Chen et al., 2022) enhances the classic Dual Contouring algorithm by refining UDFs using learned priors; however, it struggles to generalize to shapes with arbitrary topology and unseen geometric configurations. A recent example is DualMeshUDF (Zhang et al., 2023), which uses an adaptive octree and computes each mesh vertex by solving a quadratic error function (QEF) that fits the local unsigned field (analogous to finding the intersection of estimated tangent planes). These approaches can directly place vertices on the surface without requiring sign flips, enabling it to capture sharp edges and even non-manifold structures that Marching Cubes variants might miss. However, ensuring topologically correct connectivity becomes challenging for such dual methods, as they all rely on the assumption that the solution space will not be far from the grid cell and then minimizes towards a point inside of the cell close to the solution. The erroneous connectivity appears when the reconstructed surface point computed for an octree cell lies outside the cell. If these noisy points are discarded or clamped, the mesh quality is compromised by holes (see Figure 3 bottom row). In practice, DualMeshUDF filters out vertices that fall outside their cell, which means those cells end up with no mesh vertex and thereby produce a hole in the reconstructed mesh. This situation tends to occur when a non-empty cell covers surface regions on both sides of its medial axis, causing the QEF solution to drift out of bounds. One might refine the octree further to alleviate this issue, but increasing resolution explosively raises the mesh complexity. For example, in our experiments the Klein bottle model (see Figure 10) required over $13k$ vertices with DualMeshUDF to avoid holes, whereas a far coarser mesh (roughly $3.6k$ vertices) suffices with our method. Moreover, DualMeshUDF and similar octree methods impose a maximum depth limit, so they cannot indefinitely subdivide to resolve every ambiguous cell. Hence, some small holes or missed facets may remain if the depth limit is reached.
3. Method
The input to our method is a continuous UDF that provides both unsigned distance values $d=F(\mathbf{x})$ given a point $\mathbf{x}\in\mathcal{R}^{3}$, which measures proximity to the unknown iso-surface $\chi$, and gradients $\mathbf{n}(\mathbf{x})=\nabla F(\mathbf{x})/\|\nabla F(\mathbf{x})\|$, which indicate the direction of value increase. In this section, we introduce an iterative optimization algorithm to compute a high-quality triangle mesh $\mathcal{M}=(\mathcal{X},\mathcal{F})$ that approximates $\chi$ while faithfully capturing non-manifold structures, open boundaries, and sharp features.
Unlike grid-based methods that subdivide space using a fixed structure (e.g., an adaptive octree) to locate cells intersecting $\chi$, lacking the flexibility to capture thin structures and often misses features at low resolutions (see Figure 4), our method begins by initializing $N$ movable seeds with gradients $\{\mathbf{x}_{i},\mathbf{n}_{i}\}_{i=1}^{N}$. These seeds are sampled uniformly within the bounding volume and projected onto the implicit iso-surface $\chi$ using their corresponding gradients. We then partition the domain into regions via a Voronoi decomposition induced by the seed set $\{\Omega_{i}\}_{i=1}^{N}$.
The surface reconstruction is formulated as an iterative optimization of two energy terms: a local tangent energy and a global repulsive energy. The local $L_{1}$ tangent energy adjusts the position of each seed $\mathbf{x}_{i}$ within its corresponding Voronoi cell $\Omega_{i}$ to better fit the local geometry with features. The global repulsive energy redistributes the seeds to improve sampling uniformity on $\chi$ and prevent crowding, which enhances numerical stability and triangle quality in the final mesh. After convergence, the output triangle mesh is extracted as the dual of the geodesic Voronoi diagram of the optimized seeds. The full algorithm is summarized in Alg. 1.
Figure 4. Dual-contouring-based method (Zhang et al., 2023) uses voxel grids struggles to reconstruct UDF iso-surfaces with thin layers, often resulting in zig-zag connectivity and holes when using a relatively small number of vertices. In contrast, our method, based on a Voronoi partition, avoids these issues and produces clean, consistent connectivity.
Figure 5. Illustration of why our formulation uses $L_{1}$ tangent energy with Voronoi partitions, rather than $L_{2}$ tangent energy with voxel grids.
3.1. $L_{1}$ Tangent Energy
We optimize each seed’s position $\mathbf{x}_{i}$ based on the UDF values and gradients of $M$ points $\{\mathbf{p}_{i_{m}}\}_{m=1}^{M}$ sampled within its corresponding Voronoi cell $\Omega_{i}$ (we set $M=100$ in our experiments). Intuitively, the updated seed position should be consistent with the UDF value $d_{i_{m}}$ and the gradient direction $\mathbf{n}_{i_{m}}$ at each sample $\mathbf{p}_{i_{m}}$. Each sample defines an estimated tangent plane to the iso-surface $\chi$, given by the relation $\mathbf{n}_{i_{m}}^{\top}(\mathbf{p}_{i_{m}}-\mathbf{x}_{i})-d_{i_{m}}=0$. One may enforce this intuition by minimizing the deviation of the seed $\mathbf{x}_{i}$ from all such tangent planes using a quadratic energy, as proposed by Zhang et al. (2023) in the context of voxel grids:
(1) $\mathbf{x}_{i}^{\star}=\operatorname*{\arg\!\min}_{\mathbf{x}_{i}}\sum_{m=1}^{M}\Big\lVert\mathbf{n}_{i_{m}}^{\top}(\mathbf{p}_{i_{m}}-\mathbf{x}_{i})-d_{i_{m}}\Big\rVert_{2}^{2}.$
This approach, however, suffers from a longstanding limitation: the optimized point may drift outside its associated cell (Ju et al., 2002; Schaefer and Warren, 2002), resulting in invalid connectivity (see Figure 5 (a)). While enforcing hard cell-boundary constraints can prevent this drift, it often pulls the solution away from the actual iso-surface, as illustrated in Figure 5 (b).
To overcome these issues, we introduce a constrained $L_{1}$-norm tangent energy formulated over a Voronoi-based partition. Our $L_{1}$ tangent energy for seed $\mathbf{x}_{i}$ is defined as:
(2) $E_{tan}(\mathbf{x}_{i})=\sum_{m=1}^{M}\Big\lVert\mathbf{n}_{i_{m}}^{\top}(\mathbf{p}_{i_{m}}-\mathbf{x}_{i})-d_{i_{m}}\Big\rVert_{1},$
summing over all samples in $\Omega_{i}$. The optimal seed position (approximating the surface point closest to all these tangent planes) is found by minimizing this energy subject to the constraint that $\mathbf{x}_{i}$ stays within its Voronoi cell boundaries:
[…]
where the second term in the objective is a proximity regularizer that encourages each seed to remain close to its initial position $\mathbf{x}_{i}^{0}$, preventing excessive displacement during optimization (we set $\gamma=10$ in our experiment). Note that minimizing Equation 2 is equivalent to solving the following $L_{1}$ linear program:
(4) $\arg\min_{\mathbf{x}_{i}}\,\|A\,\mathbf{x}_{i}-b\|_{1},$
where $A$ is a $3\times 3$ matrix and $b$ is a $3\times 1$ vector (3 is the UDF dimension).
Figure 6. Our method successfully reconstructs non-manifold structures from UDFs, whereas the double-covering-based method (Hou et al., 2023) returns a two-layered mesh around non-manifold regions.
This constrained $L_{1}$ formulation ensures that $\mathbf{x}_{i}$ aligns with the local geometry of the zero-level set as inferred from samples within its Voronoi cell, while staying strictly within its designated region. This improves surface fidelity and avoids violations at cell boundaries. Moreover, when $\Omega_{i}$ overlaps a non-manifold or sharp feature, the optimization naturally pulls $\mathbf{x}_{i}$ onto that feature, ensuring it is accurately preserved in the final mesh.
In contrast, as shown in Figure 6, double-covering-based methods (Hou et al., 2023; Chen et al., 2025) instead return duplicated surface layers around non-manifold regions and cannot recover true connectivity.
3.2. Feature-aware Repulsive Energy
To prevent numerical instability caused by seed crowding and to improve triangle quality in the final mesh, we optimize the spatial distribution of seed positions $\{\mathbf{x}_{i}\}_{i=1}^{N}$. This can be achieved using a particle-based repulsion model (Witkin and Heckbert, 1994; Zhong et al., 2013; Wang et al., 2025) optimized with L-BFGS. The inter-particle repulsive energy between each pair of distinct seeds $(i,j)$ is defined as:
(5) $E_{rep}=\sum_{i}\sum_{j\neq i}E_{rep}^{ij}=\sum_{i}\sum_{j\neq i}\exp\!\Big(-\frac{\|\mathbf{x}_{j}-\mathbf{x}_{i}\|^{2}}{2\,\sigma^{2}}\Big),$
where $\sigma$ (the kernel width) is the standard deviation of the Gaussian kernel. We set $\sigma$ at each optimization step to the minimum Euclidean distance between any pair of seeds (see Alg. 1, line 4).
However, allowing seeds to move freely within the space without considering geometric features can lead to undesirable effects. For example, seeds may drift completely away from the iso-surface $\chi$ during optimization. Inspired by Wang et al. (2025), we introduce a feature-aware gradient projection strategy to mitigate this issue using the geometric information extracted from the $L_{1}$ tangent energy defined in Sec. 3.1. More specifically, We examine the null space of $A$ in Equation 4 via singular value decomposition (SVD), writing $A=U\,\Sigma\,V^{T}$. If $\text{rank}(A)=r$, then the last $(\dim-r)$ columns of $V$ (denoted $V_{0}$) form an orthonormal basis for the null space of $A$. Our strategy is to project the force $\mathbf{g}^{i}=\sum_{j}\nabla E_{rep}^{ij}$ onto this null space:
(6) $\mathbf{g}_{proj}^{i}=V_{0}\,V_{0}^{\top}\,\mathbf{g}^{i},$
ensuring that the gradient update lies in the feature-preserving subspace. In practice, this restricts movement to feature-consistent directions. For example, a seed that has converged to a non-manifold edge will only be updated along that edge, preserving sharp or singular structures during optimization.
In summary, our method produces high-quality triangle meshes, as shown in Figure 7, due to the repulsive energy formulation combined with feature-aware gradient projection.
Figure 7. Our method reconstructs the UDF iso-surface with better triangle quality compared to the marching-cubes-based approach (Ren et al., 2023).
3.3. Mesh Reconstruction via the Dual of a Geodesic Voronoi Diagram
After optimizing the seeds, the next step is to connect them to form a triangle mesh. A natural approach is to use the dual of the Voronoi diagram (i.e., the Delaunay triangulation) to establish connectivity among seeds. However, a standard Voronoi diagram computed in the entire domain is unrestricted; it may erroneously connect seeds on different layers of the iso-surface even if they are far apart on the surface. Even a restricted Voronoi diagram can fail in thin-plate scenarios: a single site may dominate multiple surface regions when the iso-surface is a thin plate and the number of seed points is insufficient. For example Figure 8 (a), the dual of the VD contains extra edge from $\mathbf{x}_{A}$ to $\mathbf{x}_{B}$ as their Voronoi cells are adjacent.
To address these issues, we propose a reconstruction method based on the Geodesic Voronoi Diagram (GVD) defined on the iso-surface itself, shown in Figure 8 (b). Instead of computing the Voronoi diagram within the bounding box using the Euclidean distance $d_{\text{voro}}$, we compute a GVD on the zero-level set using the geodesic distance $d_{\text{geo}}$.
Computing an exact GVD on an implicit zero-level set (without an explicit mesh) is challenging, so we develop a sampling-based approximation. We first densely sample points in space then project on iso-surface $\chi$ using their UDF values and gradients, and construct a $k$-nearest neighbor (KNN) graph on these samples (using $k=20$ in our experiments). Next, we prune edges from this graph that likely do not lie on the surface. In particular, we remove any edge if either (1) the unsigned distance value at its midpoint is above a threshold (i.e., $F\big(1/2(\mathbf{x}_{i}+\mathbf{x}_{j})\big)>\epsilon_{udf}$), indicating the midpoint is off the surface, or (2) the UDF gradient at either endpoint is too aligned with the edge’s direction (quantified by the dot product exceeding $\epsilon_{grad}$), indicating the edge is not tangent to the surface. After this filtering, we treat each seed as a source and run a sweeping algorithm (Xin et al., 2022) on the remaining graph to label every sample with its nearest seed in terms of geodesic distance, effectively simulating a flooding process on the surface. Finally, we construct the triangle connectivity (the dual of the approximate GVD) by examining the labeled samples: for each surface sample labeled $i$ (nearest to seed $\mathbf{x}_{i}$), if it has neighboring samples labeled $j$ and $k$ (with $j\neq i$ and $k\neq i$), we introduce a triangle connecting seeds $(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{x}_{k})$ in the mesh. Each such triangle corresponds to a junction where three Voronoi regions meet on the surface.
3.4. Mesh Thinning
The initial mesh produced by the geodesic Voronoi diagram (GVD) reconstruction is inherently thick, as the dual of the GVD introduces flat but solid volumetric cells. As a result, the extracted surface is not yet a thin manifold. To eliminate these artifacts and recover a clean surface, we apply a thinning procedure inspired by (Wang et al., 2022, 2024, 2025).
We begin by detecting all solid tetrahedral cells embedded in the triangle mesh and iteratively prune triangle pairs from each tetrahedron without altering the overall mesh topology. Figure 9 illustrates the thinning operation on a single tetrahedron. Repeatedly removing such pairs from all detected tetrahedra collapses the flat volumes, yielding a thinner and more consistent surface mesh.
Figure 8. Illustration of constructing an approximate geodesic Voronoi diagram (GVD) from a standard Voronoi diagram. (a) Points are densely sampled in space and around the seeds, then projected onto the zero-level set. (b) The approximate GVD is computed by running a sweeping algorithm from each seed. See Sec. 3.3 for details.
However, not all volumetric artifacts are tetrahedral; in practice, some remaining solid cells form more complex polyhedral structures. To address these cases, we iteratively remove small manifold components (defined as connected components with fewer than 10 faces in our experiments) until none remain. This post-processing step eliminates residual non-manifold polyhedral artifacts, resulting in a cleaner, more compact, and topologically valid surface mesh.
The effectiveness of our results relies on both the optimization and the thinning process. As shown in Figure 10, we conduct an ablation study in which our thinning procedure is applied as a post-processing step to both the output of DualMeshUDF (Zhang et al., 2023) and our own method.
Figure 9. Illustration of the thinning process for a flat tetrahedron $t_{1}$ in a 3D reconstructed mesh. (a) A triangle pair ($f_{1}$, $f_{2}$) sharing a manifold edge with highest UDF value on its centroid is selected for pruning. (b) After removing this pair, the remaining faces ($f_{3}$, $f_{4}$) constitute the pruned surface.
Figure 10. We apply our thinning method described in Sec. 3.4 to both the dual-contouring-based method (Zhang et al., 2023) and our approach. The comparison illustrates that the high-quality of our results relies on both the optimization and the thinning process. Non-manifold regions are highlighted in red.
Data: Unsigned distance function $F(\mathbf{x})$ with gradients
Result: Reconstructed mesh $\mathcal{M}_{r}$
1 Initialize NN seeds $\{\mathbf{x}_{i}\}_{i_{1}}^{N}$
2 while optimization not converged do
3   Update seeds using $L_{1}$ tangent energy in Sec. 3.1;
4   Recompute the repulsive kernel width $\sigma$;
5 while stopping condition not satisfied do
6      Project seeds $\{\mathbf{x}_{i}\}$ using $F(\mathbf{x}_{i})$ and its gradients;
7      Re-compute Voronoi diagram;
8      Sample each Voronoi cell of $\{\mathbf{x}_{i}\}_{i=1}^{N}$;
9    for each seed $\mathbf{x}_{i}$ do
10       for each neighbors $\mathbf{x}_{j}$ of $\mathbf{x}_{i}$ do
11            Compute $E_{rep}^{ij}$ and $\mathbf{g}^{ij}$;
12         end for
13        Sum the total force $\mathbf{g}^{i}=\sum_{j}\mathbf{g}^{ij}$;
14         Compute projected $\mathbf{g}_{proj}^{i}$ using Eq. 6;
15      end for
16     Sum the total energy $E_{rep}$ using Eq. 5;
17      Run L-BFGS with $E_{rep}$ and $\mathbf{g}_{proj}=\sum_{i}\mathbf{g}_{proj}^{i}$;
18
19   end while
20  Project seeds $\{\mathbf{x}_{i}\}$ using $F(\mathbf{x}_{i})$ and its gradients;
21
22 end while
23Compute reconstructed mesh in Sec. 3.3;
Apply mesh thinning in Sec. 3.4;
ALGORITHM 1 Optimization
4. Experiments
We implemented VoroUDF in Python with C++ bindings and conducted all experiments on a MacBook Pro equipped with an Apple M3 Pro chip and 36GB of memory. All baseline methods were executed on a Linux machine with a 16-core AMD Ryzen processor and an NVIDIA RTX 3090 GPU. All 3D models in the figures are rendered using BlenderToolbox (Liu, 2018). We will release our code on GitHub publicly upon acceptance.
Figure 11. Ablation study showing the effect of the termination threshold $\mathbf{\delta}$ on runtime and reconstruction quality. Lower values of $\mathbf{\delta}$ reduce the edge Chamfer error but incur higher computational cost. We choose $\mathbf{\delta}=10^{-3}$ as a practical trade-off between accuracy and efficiency.
Optimization
The full optimization procedure is summarized in Alg. 1, with a termination threshold denoted by $\mathbf{\delta}$. The inner loop (line 5) terminates when either the relative change in the L-BFGS gradients falls below $\mathbf{\delta}$ or the maximum number of iterations is reached. The outer loop (line 2) stops when the inner loop finishes without performing any updates, indicating convergence. We present an ablation study in Figure 11 to illustrate how varying $\mathbf{\delta}$ affects both runtime and reconstruction quality. In our experiments, we use $\mathbf{\delta}=10^{-3}$ as it achieves low edge Chamfer error while maintaining reasonable runtime.
Comparison Methods
We compare our method against seven baseline approaches, which fall into three main categories. The marching-cubes-based methods include MeshUDF (Guillard et al., 2022), CapUDF (Zhou et al., 2022), GeoUDF (Ren et al., 2023), and the learning-based method NSDUDF (Stella et al., 2024). For MeshUDF, CapUDF, and GeoUDF, we use the standalone mesh extraction modules provided in their released code. Since CapUDF’s extraction duplicates vertices per grid cell, we apply the same post-processing step used in MeshUDF to remove duplicate vertices for consistency. For NSDUDF, we evaluate the publicly released model trained on the ABC dataset across all benchmarks. We also compare against two double-covering-based methods, DCUDF (Hou et al., 2023) and DCUDF2 (Chen et al., 2025), and one dual-contouring-based method, DualMeshUDF (Zhang et al., 2023). All methods are evaluated under comparable reconstruction settings, with a similar number of vertices (#v\#v) in the output mesh. Additional comparisons at higher resolutions are included in the Supplementary Material.
Evaluation Metrics
To evaluate the quality of the reconstructed meshes, we adopt several standard metrics. For mesh regularity, we report Triangle Quality (TQ(Frey and Borouchaki, 1999) and Topology Error (TD). To assess geometric fidelity between the reconstructed surface and the ground truth, we use the $L_{1}$ Chamfer Error (CD) and the Hausdorff Error (HD). To evaluate sharp feature reconstruction, we report the $L_{1}$ Edge Chamfer Error (ECD(Chen and Zhang, 2021). For non-manifold regions, we introduce the Non-Manifold Chamfer Error (NM-CD) to measure reconstruction accuracy. Due to space constraints, detailed definitions of TQ, $L_{1}$-CD, $L_{1}$-ECD, and HD are provided in the Supplementary Material.
Topology Error (TD)
We use the Euler characteristic as a measure of mesh topology and report the absolute difference between the Euler characteristic of the source UDF isosurface and that of the reconstructed mesh. This metric reflects the topological discrepancy between the reconstruction and the ground truth.
Non-Manifold Chamfer Error (NM-CD)
We detect non-manifold edges (shared by more than two faces) and non-manifold vertices in both the source and reconstructed meshes. We then densely sample points on these regions and compute the bidirectional Chamfer distance between the two sets of samples. The $L_{1}$ error of this distance is reported as the NM-CD metric, quantifying the geometric accuracy of reconstructed non-manifold structures.
4.1. Comparisons on Non-Manifold Models
To quantitatively evaluate our method on non-manifold geometry, we construct a set of 10 synthetic non-manifold models (see inset of our results). Each model is generated by selecting a mesh from the ABC dataset (Koch et al., 2019), cutting it with a plane, and then merging the two cut parts with a narrow overlapping band (see Figure 12, source). We densely sample points on each of the two cut meshes and identify those near the opposing part, which we highlight in red in Figure 12. These red samples approximate the ground-truth non-manifold edges in the merged mesh. Quantitative results are reported in Table 4, and corresponding qualitative comparisons are shown in Table 4. We select one representative method from each category to compare in the figure. More results are provided in the Supplementary Material. Among all methods evaluated, ours is the only one that faithfully reconstructs non-manifold structures. It is worth noting that the double-covering-based methods (Hou et al., 2023; Chen et al., 2025) completely fail to recover non-manifold structures, as they produce double-layered meshes by design, resulting their metric $\textit{NM-CD}=\text{Inf}$.
Table 1. Quantitative comparison on 10 synthetic models with non-manifold edges against seven state-of-the-art methods, using a similar number of vertices (#v\#v) in the reconstructed meshes. The best scores are shown in bold, and the second best scores are underlined.
[…]
MeshUDF 7.5k 64.09 33.07 32.36 0.65 157.2 1017.12
CapUDF 6.3k 53.25 126.86 25.91 0.57 261.3 Inf
GeoUDF 4.5k 22.43 36.61 37.28 0.65 142.1 1499.15
NSDUDF 3.6k 20.29 22.08 99.55 0.64 84.7 267.09
DCUDF 8.2k 80.96 32.27 100.29 0.71 45.1 Inf
DCUDF2 9.1k 64.76 28.13 110.31 0.72 45.0 Inf
DualMeshUDF 5k 17.89 11.35 4.92 0.52 744.3 375.24
Ours 2.9k 17.04 6.74 4.05 0.84 43.6 14.35
4.2. Comparisons on CAD models
We present quantitative comparison results on 100 CAD models taken from Xu et al. (2024) in Tab. 2, and corresponding qualitative comparisons in Figure 14. Our method consistently outperforms all baselines across nearly all metrics, while maintaining a comparable number of vertices in the reconstructed mesh. Additional results using denser resolutions for the seven baseline methods are provided in the Supplementary Material.
Table 2. Quantitative comparison on the 100 ABC models (Xu et al., 2024) against seven state-of-the-art methods, using a similar number of vertices (#v\#v) in the reconstructed meshes. The best scores are shown in bold, and the second best scores are underlined.
[…]
MeshUDF 4.2k 58.80 28.39 42.78 0.65 47.16
CapUDF 6.7k 44.10 94.71 29.36 0.52 189.64
GeoUDF 2.6k 19.59 31.58 52.33 0.67 59.07
NSDUDF 2.3k 17.17 15.78 103.82 0.65 13.5
DCUDF 4.8k 91.12 51.34 123.44 0.62 1.17
DCUDF2 3.8k 109.57 53.51 130.63 0.65 1.44
DualMeshUDF 3.1k 30.62 10.60 4.98 0.52 694.74
Ours 2.2k 13.07 5.85 2.38 0.83 0.28
4.3. Comparisons on Garment Models
We further evaluate our method against seven state-of-the-art baselines on 100 garment models from the DeepFashion3D dataset (Zhu et al., 2020). Quantitative results are reported in Tab. 3, and qualitative comparisons are shown in Figure 14. Our method would introduces non-manifold edges (highlighted in red) in regions under the armpit, where two layers of the surface nearly coincide. This contributes to a slightly higher topology error (TD) compared to double-covering-based methods (Hou et al., 2023; Chen et al., 2025), which avoid such configurations by design.
Table 3. Quantitative comparison on the top 100 DeepFashion3D (Zhu et al., 2020) models against seven state-of-the-art methods, using a similar number of vertices (#v\#v) in the reconstructed meshes. The best scores are shown in bold, and the second best scores are underlined.
[…]
MeshUDF 3.9k 64.7 31.3 0.61 75.68
CapUDF 5.2k 27.39 30.90 0.53 322.11
GeoUDF 2.6k 16.79 31.77 0.61 76.13
DCUDF 3.8k 107.36 78.15 0.70 1.23
DCUDF2 4.1k 89.86 68.85 0.71 1.18
DualMeshUDF 3k 14.17 27.01 0.52 499.48
Ours 2.6k 13.09 20.65 0.74 1.76
5. Limitations & Conclusion
We have introduced VoroUDF, a surface reconstruction algorithm that generates high-quality triangle meshes from unsigned distance fields. Unlike all prior work on this task, our results preserve faithful non-manifold structures, sharp features, and open surfaces. Our method leverages a Voronoi-based partitioning framework, combined with an $L_{1}$ tangent energy and a feature-aware repulsive energy, to achieve state-of-the-art performance across a wide range of input scenarios.
Our Voronoi-based spatial partitioning eliminates the need for subdividing voxel grids and is better suited for preserving thin structures. However, it requires seed points to be initialized near such features; without adequate initialization, small-scale geometry may be missed or poorly reconstructed. In our current implementation, seeds are uniformly sampled across the domain, which leaves coverage to chance in complex regions. In future work, we plan to explore adaptive or data-driven seed initialization strategies to ensure sufficient sampling around thin or intricate structures.
Our method demonstrates strong empirical performance in preserving complex topologies such as non-manifold structures, it unfortunately lacks a formal theoretical guarantee of topological correctness. In some cases, our reconstruction may introduce very small amount of non-manifold edges (comparing to dual-countring-based
method (Zhang et al., 2023)) even when the input source is a true manifold (see inset). This behavior is most evident in regions where surface layers nearly coincide, such as the armpit area in garment models. Additionally, some triangles near open boundaries may be erroneously removed due to over-thinning. As future work, we plan to incorporate topological validation or correction mechanisms to improve robustness in these challenging cases.
Our $L_{1}$ tangent energy and repulsive energy formulations are designed to operate reliably on clean UDFs with accurate gradients. However, the effect of input noise, either in the distance values or gradients, has not been systematically studied. Investigating the robustness of our method under noisy conditions and exploring denoising techniques or regularization strategies will be an important direction for future research.
Our work focuses on reconstructing UDFs with complex geometries and topologies. With it, we aim to contribute to a growing body of research that helps bridge the gap between the flexible modeling practices of 3D artists and the geometric quality and fidelity requirements of downstream applications. We also seek to help connect the 3D generative deep learning and geometry processing communities, enabling generated shapes to be used in real-world tasks like rendering, physical simulation, and animation.
Acknowledgements.
The Geometry and the City lab at Columbia University is supported by generous gifts from nTop, Adobe and Braid Technologies.
Appendix A Supplementary Material
Appendix B Evaluation Metrics
Triangle Quality (TQ)
To evaluate triangle quality, we use the TQ indicator (Frey and Borouchaki, 1999), where a value closer to $1.0$ indicates a triangle that is nearly equilateral. We use the following equation to evaluate the triangle quality,
(7) $\textit{TQ}(t)=\frac{6}{\sqrt{3}}\cdot\frac{S_{t}}{p_{t}\cdot h_{t}}$
where $S_{t}$, $p_{t}$, and $h_{t}$ denote the area, half-perimeter, and the longest edge length of the triangle $t$, respectively.
$L_{1}$ Chamfer Error (CD)
The Chamfer error measures the overall geometric discrepancy between the reconstructed surface mesh and the ground truth shape. In our experiments, we report the $L_{1}$ version of this metric to evaluate reconstruction accuracy.
$L_{1}$ Edge Chamfer Error (ECD)
To assess the accuracy of sharp feature reconstruction, we adopt the edge Chamfer distance introduced by Chen and Zhang (2021), computed using the $L_{1}$ norm.
Hausdorff Error (HD)
To measure the difference between the reconstructed mesh and the source, we use the Hausdorff Error as the indicator. Here, $\textit{HD}^{1}$ is the one-sided Hausdorff distance from the original surface to the surface reconstructed from MAT, and $\textit{HD}^{2}$ is the distance in reverse side. All Hausdorff distances are evaluated as percentages of the distance over the diagonal lengths of the models’ bounding box. The HD is the maximum of $\textit{HD}^{1}$ and $\textit{HD}^{2}$. The Hausdorff distance can reveal if there is anything missing or redundant in the reconstructed geometry (e.g., a hole or a floating piece).
B.1. More Comparisons on Non-Manifold Models
We compare our method against seven state-of-the-art baselines on 10 synthetic models containing non-manifold edges, evaluated at higher resolution ($2^{7}$). Quantitative results are reported in Table 4, and qualitative comparisons are shown in Figure 15.
Table 4. Quantitative comparison on 10 synthetic models with non-manifold edges, evaluated against seven state-of-the-art methods with higher resolution ($2^{7}$). The best scores are shown in bold, and the second best scores are underlined.
[…]
MeshUDF 121k 23.64 7.50 31.47 0.66 479.7 1267.02
CapUDF 103k 18.56 9.30 20.96 0.58 1531.7 Inf
GeoUDF 71k 17.12 8.55 32.52 0.67 521.9 843.62
NSDUDF 60k 17.02 4.06 58.08 0.64 257.3 784.42
DCUDF 159k 27.41 6.78 124.62 0.71 45.5 Inf
DCUDF2 175k 21.72 5.80 120.64 0.72 45.4 Inf
DualMeshUDF 8.4k 17.21 1.35 3.78 0.61 22k 418.01
Ours 2.9k 17.04 6.74 4.05 0.84 43.6 14.35
Figure 15. Qualitative comparison on 10 synthetic models with non-manifold edges, evaluated against seven state-of-the-art methods at higher resolution ($2^{7}$). Non-manifold edges are highlighted in red.
B.2. More Comparisons on CAD models
We compare our method against seven state-of-the-art baselines on 100 CAD models (Xu et al., 2024), evaluated at higher resolution ($2^{7}$). Quantitative results are presented in Table 5, and qualitative comparisons are shown in Figure 16 and Figure 17.
Table 5. Quantitative comparison on 100 ABC models (Xu et al., 2024), evaluated against seven state-of-the-art methods with higher resolution ($2^{7}$). The best scores are shown in bold, and the second best scores are underlined.
[…]
MeshUDF 70k 33.5 6.3 43.35 0.67 239.15
CapUDF 111k 14.90 6.16 18.42 0.53 813.06
GeoUDF 42k 13.13 6.59 38.59 0.68 296.63
NSDUDF 36k 13.04 2.79 62.83 0.66 94.02
DCUDF 67k 41.24 12.32 137.10 0.64 0.32
DCUDF2 95k 20.69 7.98 96.81 0.60 0.35
DualMeshUDF 48k 29.19 0.98 2.83 0.61 10053.07
Ours 2.2k 13.07 5.85 2.38 0.83 0.28
Figure 16. Qualitative comparison on 100 ABC models (part 1), evaluated against seven state-of-the-art methods at higher resolution ($2^{7}$). Non-manifold edges are highlighted in red.
Figure 17. Qualitative comparison on 100 ABC models (part 2), evaluated against seven state-of-the-art methods at higher resolution ($2^{7}$). Non-manifold edges are highlighted in red.
B.3. More Comparisons on garment models
We further compare our method with seven state-of-the-art methods on (first) $100$ garment shapes from DeepFashion3D (Zhu et al., 2020) dataset. The results from seven state-of-the-art methods are generated at higher resolution ($2^{7}$), shown in Table 6, Figure 18 and Figure 19.
Table 6. Quantitative comparison on first 100 DeepFashion3D models (Zhu et al., 2020), evaluated against seven state-of-the-art methods with higher resolution ($2^{7}$). The best scores are shown in bold, and the second best scores are underlined.
[…]
MeshUDF 67k 20.53 8.20 0.61 726.00
CapUDF 77k 13.36 9.05 0.53 2055.21
GeoUDF 37k 11.95 10.07 0.62 237.34
DCUDF 53k 29.36 37.39 0.72 4.43
DCUDF2 58k 20.04 35.73 0.71 3.33
DualMeshUDF 51k 12.08 7.62 0.58 8131.65
Ours 2.6k 13.09 20.65 0.74 1.76
Figure 18. Qualitative comparison on 100 DeepFashion3D models (part 1), evaluated against seven state-of-the-art methods at higher resolution ($2^{7}$). Non-manifold edges are highlighted in red.
Figure 19. Qualitative comparison on 100 DeepFashion3D models (part 2), evaluated against seven state-of-the-art methods at higher resolution ($2^{7}$). Non-manifold edges are highlighted in red.

### v2

@@
 Critically, it does so without resorting to an excessively dense discretization, producing lightweight meshes that are directly portable to realtime and interactive applications.
-2. Related Work
 Figure 3.
 Illustration of why marching-cubes-based methods (top row), double-covering-based methods (middle row), and dual-contouring-based methods (bottom row) fail to preserve the non-manifoldness of the iso-surface.
+2. Related Work
 UDFs can represent surfaces of arbitrary topology, making them particularly valuable in domains such as artistic modeling, manufacturing, and industrial design, where accurate recovery of complex surface structures is critical.
@@
 In this section, we introduce an iterative optimization algorithm to compute a high-quality triangle mesh $\mathcal{M}=(\mathcal{X},\mathcal{F})$ that approximates $\chi$ while faithfully capturing non-manifold structures, open boundaries, and sharp features.
-Unlike grid-based methods that subdivide space using a fixed structure (e.g., an adaptive octree) to locate cells intersecting $\chi$, lacking the flexibility to capture thin structures and often misses features at low resolutions (see Figure 4), our method begins by initializing $N$ movable seeds with gradients $\{\mathbf{x}_{i},\mathbf{n}_{i}\}_{i=1}^{N}$.
+Unlike grid-based methods that subdivide space using a fixed structure (e.g., an adaptive octree) to locate cells intersecting $\chi$, lacking the flexibility to capture thin structures and often misses features at low resolutions (see Figure 4), our method begins by initializing NN movable seeds with gradients $\{\mathbf{x}_{i},\mathbf{n}_{i}\}_{i=1}^{N}$.
 These seeds are sampled uniformly within the bounding volume and projected onto the implicit iso-surface $\chi$ using their corresponding gradients.
@@
 (4) $\arg\min_{\mathbf{x}_{i}}\,\|A\,\mathbf{x}_{i}-b\|_{1},$
-where $A$ is a $3\times 3$ matrix and $b$ is a $3\times 1$ vector (3 is the UDF dimension).
+where AA is a $3\times 3$ matrix and $b$ is a $3\times 1$ vector (3 is the UDF dimension).
 Figure 6. Our method successfully reconstructs non-manifold structures from UDFs, whereas the double-covering-based method (Hou et al., 2023) returns a two-layered mesh around non-manifold regions.
@@
 3.1.
-More specifically, We examine the null space of $A$ in Equation 4 via singular value decomposition (SVD), writing $A=U\,\Sigma\,V^{T}$.
-If $\text{rank}(A)=r$, then the last $(\dim-r)$ columns of $V$ (denoted $V_{0}$) form an orthonormal basis for the null space of $A$.
+More specifically, We examine the null space of AA in Equation 4 via singular value decomposition (SVD), writing $A=U\,\Sigma\,V^{T}$.
+If $\text{rank}(A)=r$, then the last $(\dim-r)$ columns of $V$ (denoted $V_{0}$) form an orthonormal basis for the null space of AA.
 Our strategy is to project the force $\mathbf{g}^{i}=\sum_{j}\nabla E_{rep}^{ij}$ onto this null space:
@@
 Figure 7. Our method reconstructs the UDF iso-surface with better triangle quality compared to the marching-cubes-based approach (Ren et al., 2023).
-3.3. Mesh Reconstruction via the Dual of a Geodesic Voronoi Diagram
+$3.$ Mesh Reconstruction via the Dual of a Geodesic Voronoi Diagram
 After optimizing the seeds, the next step is to connect them to form a triangle mesh.
@@
 Instead of computing the Voronoi diagram within the bounding box using the Euclidean distance $d_{\text{voro}}$, we compute a GVD on the zero-level set using the geodesic distance $d_{\text{geo}}$.
+Figure 8.
+Illustration of constructing an approximate geodesic Voronoi diagram (GVD) from a standard Voronoi diagram.
+(a) Points are densely sampled in space and around the seeds, then projected onto the zero-level set.
+(b) The approximate GVD is computed by running a sweeping algorithm from each seed.
+See Sec.
+3.3 for details.
 Computing an exact GVD on an implicit zero-level set (without an explicit mesh) is challenging, so we develop a sampling-based approximation.
@@
 Repeatedly removing such pairs from all detected tetrahedra collapses the flat volumes, yielding a thinner and more consistent surface mesh.
-Figure 8.
-Illustration of constructing an approximate geodesic Voronoi diagram (GVD) from a standard Voronoi diagram.
-(a) Points are densely sampled in space and around the seeds, then projected onto the zero-level set.
-(b) The approximate GVD is computed by running a sweeping algorithm from each seed.
-See Sec.
-3.3 for details.
 However, not all volumetric artifacts are tetrahedral; in practice, some remaining solid cells form more complex polyhedral structures.
@@
 B.3. More Comparisons on garment models
-We further compare our method with seven state-of-the-art methods on (first) $100$ garment shapes from DeepFashion3D (Zhu et al., 2020) dataset.
+We further compare our method with seven state-of-the-art methods on (first) 100100 garment shapes from DeepFashion3D (Zhu et al., 2020) dataset.
 The results from seven state-of-the-art methods are generated at higher resolution ($2^{7}$), shown in Table 6, Figure 18 and Figure 19.